OpenStudy (anonymous):

A student records the absorption spectrum for an intensely colored substance and finds a value of lamba_max =520nm. The absorbance of this solution is 0.712 at 520nm. The solution was measured in a cuvette with a path length of 1cm. (a) If the molar absorptivity is 3.56 * 10^3L/mol cm, what is the concentration of this solution? (b) what is the percent transmittance? (c) what color is absorbed by the solution? (d) what is the color of the solution?...

thomaster (thomaster):

(a) Use the lambert beer formula A = a x b x c (Absorption = extinction coefficient * path length * concentration) c = A / ( a x b ) (b) A= -log T so T = 10^-A %T = T * 100 (c) every color except the color of the solution. (d) 495–570 nm = yellow

thomaster (thomaster):

for question a, the given molar absorptivity is the extinction coefficient.